OpenStudy (anonymous):

Divide f(x) by d(x), and write a summary statement in the form indicated. f(x) = x^4 + 4x^3 + 6x^2 + 4x + 5; d(x) = x^2 + 1

OpenStudy (anonymous):

Any idea how to start?

OpenStudy (anonymous):

if i divide it gonna be x^2+4x+5

OpenStudy (anonymous):

Okay, good. So what are you unsure of? How to write the summary statement?

OpenStudy (anonymous):

I would need to see the form they indicate, but it is probably something like: ____________ divided by __________ = __________ with remainder _____

OpenStudy (anonymous):

these are the alternatives: f(x) = (x2 + 1)( x2 + 4x + 5) + 12x - 15 f(x) = (x2 + 1)( x2 + 4x + 5) f(x) = (x2 + 1)( x2 - 4x + 5) f(x) = (x2 + 1)( x2 - 4x + 5) + 12x - 15

OpenStudy (anonymous):

Oooh, okay. Well, just think for a moment about how division works. Simple division problem example: \(15 \div 3 = 5\) Why? Because: 15 = 3*5 Same thing here. You started with f(x) and divided by \(x^2+1\) and got \(x^2+4x+5\). Turn that into a multiplication equation just like above.

OpenStudy (anonymous):

i got it ..thank u :)

OpenStudy (anonymous):

My pleasure =)